OpenStudy (anonymous):

how do you solve 2(x-3)^2=16

OpenStudy (lalaly):

(x-3)=x^2-6x+9 2(x^2-6x+9)=2x^2-12x+18 2x^2-12x+18=16 2x^2-12x+2=0 divide by 2 x^2-6x+1=0

OpenStudy (lalaly):

then use the formula